# Conditional Text-to-Image Generation with Reference Guidance

### v1

@@ A.6 Text Image Editing Details
We elaborate on the extension of our model for text image editing. In line with [12], we augment the input channel by incorporating additional encoded latent obtained from masked images. Specifically, we use VAE to encode an image containing masked regions and concatenate the encoded latent with the original model input. We use the same VAE used for reference image encoding to encode the masked image, and noise is not added to the resulting latent. Consequently, we obtain an augmented input $\tilde{\mathbf{x}}_{i}^{t}\in\mathcal{R}^{(3c+1)\times h\times w}$ as depicted below:
$\tilde{\mathbf{x}}_{i}^{t}=\text{concat}(\mathbf{z}_{i}^{t},\mathbf{s}_{i},\mathbf{m}_{i},\mathbf{z}_{i}^{\text{masked}}),$ (10)
@@ B.1 Ablation Study on Languages
ControNet 0.1500 0.1200 0.0850 0.0700 0.0850 0.1100 0.1500 0.0600 0.1950 0.1063 0.1200 0.1175 119.996
SD 0.0000 0.0050 0.0100 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0038 0.0000 0.0015 93.198
TextDiffuser 0.8400 0.5900 0.6250 0.2750 0.0000 0.0000 0.0000 0.0000 0.0000 0.5825 0.0000 0.3170 96.826
Ours 0.7000 0.4800 0.5600 0.5150 0.2100 0.1150 0.3800 0.1550 0.3300 0.5638 0.2380 0.4145 97.523
Ours 0.8175 0.5850 0.6050 0.5750 0.2200 0.1000 0.4800 0.1800 0.4800 0.6456 0.2920 0.4860 97.156
@@
Non-Latin Hindi 193.788 0.3125 0.2100 0.3754
Bengali 192.716 0.3135 0.1150 0.1760
Russian 188.567 0.2998 0.3800 0.5689
Thai 203.457 0.3015 0.1550 0.2759
Greek 192.215 0.3023 0.3300 0.5520
Average 97.523 0.3006 0.4145 0.5894
@@
Per language comparison. We provide OCR Accuracy for each language and compare them with the existing methods in Table A. Remarkably, our model is shown to be the most accurate, as it surpasses existing methods across all average accuracy metrics including the average of Latin, non-Latin, and the entire language set by a significant margin. Moreover, our method achieves high English accuracy comparable to that of TextDiffuser, which was tailored for English image generation. Although TextDiffuser shows high accuracy in some languages such as English and German, it shows incapability in the rest of the language set, as it achieves 0 accuracy. In contrast to TextDiffuser, our method can generate text in all languages, and this demonstrates the multi-lingual capability of our model.
Latin v.s. Non-Latin. Our model exhibits higher OCR Accuracy scores in Latin languages such as English, Italian, German, and French (Tables A and B). It is noteworthy that for Russian, Thai, and Greek, our model is exclusively trained on synthetic images, yet it achieves comparable or superior FID and CLIP scores compared to the results for Latin languages. Given that the model was pretrained on a large-scale English dataset [12], we speculate that this contributes as one factor to its superior performance in Latin languages.
Different Levels of Complexity Among Languages. We note that different languages present different levels of complexity, with some being more challenging to generate accurately (Tables A and B). For instance, the size of the alphabet in English is 26, whereas in Thai, the size is 59. In addition, in languages such as Thai and Greek, diacritics are combined with other letters, and extra complexity is added. Some of the predictions that our model makes look similar to the actual word, but are evaluated to be incorrect due to the mis-generation of diacritics. Besides, the shapes of characters in non-Latin languages (e.g., Bengali), are notably more intricate than Latin characters, whose alphabet typically requires a greater number of strokes to form.
@@
0.5 0.5 7 38.104 0.3454 0.5984
0.5 1.5 6 39.932 0.3448 0.5925
0.5 2.5 5 42.192 0.3434 0.6173
0.5 3.5 4 44.680 0.3419 0.6251
0.5 4.5 3 49.827 0.3317 0.6385
- 0 7.5 38.593 0.3454 0.5826
Table D: Evaluation results on MARIO-EVAL with varying portions of scheduled guidance $\gamma$. $\gamma=0$ denotes the result of CFG without the guidance scheduling.
@@ B.2 Ablation Study on CFG scheduling
$\begin{split}\hat{\epsilon}^{t}&=\underbrace{\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)}_{\text{Unconditional}}\\ &\underbrace{+\omega_{\text{prompt}}\Big{[}\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\mathbf{c},t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{Caption Guidance}}\\ &+\underbrace{\omega_{\text{ref}}\Big{[}\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\varnothing,t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{Reference Guidance}}\\ &+\underbrace{\omega_{\text{all}}\Big{[}\epsilon(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\mathbf{c},t)-\epsilon(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{All Guidance}},\end{split}$
@@
Impact of the Scheduling Speed. We first study the impact of CFG scheduling by varying $\rho_{\text{speed}}$, the increase/decrease rate of the guidance scale, in Table D. We observe a rapid decrease in $\omega_{\text{ref}}$ with a rapid increase in $\omega_{\text{ref}}$ result in a decrease in the OCR Accuracy and improvements in FID and CLIP scores. As $\omega_{\text{ref}}$ decreases rapidly, the impact of the prompt gets higher and the reference condition becomes less influential in the generation process. As a consequence, non-target words within the prompt gain increased influence and affect the generation, resulting in the inclusion of non-target words in the output (Figure A).
Impact of Different Portions of Scheduled Guidance. We examine the effect of different portions of scheduled guidance $\gamma$ within the total amount of guidance (Table D). The result of CFG without scheduling is also provided for comparison. We observe the general trend of a decrease in CLIP scores and an increase in OCR Accuracy as $\gamma$ increases. We speculate that the initial guidance scale of the reference condition plays an important role in establishing the initial layout of the text to be generated which, in turn, influences the effect of the prompt condition and consequently affects CLIP and FID.
@@ B.3 Ablation Study on Hyper Parameters
Impact of synthetic loss weight. We first analyze the impact of the synthetic loss weight in Table F. We note that the synthetic loss weight mainly impacts the CLIP as the text prompts of the real dataset better describe the content of an image. We observe the model trained with a synthetic loss weight of 0.5 achieves a lower CLIP than the one trained with $1.0$, which validates the effectiveness of lowering the weight of reconstruction loss when synthetic images are provided.
Impact of recognition loss weight. We analyze the impact of the recognition loss weight on accuracy and the quality of the generated results in Table F. We note that the OCR accuracy improves as the model is trained with increased recognition loss weight. In general, models with higher loss weight achieve better OCR Accuracy and a worse score on FID. We choose the recognition weight as $0.025$ for a proper trade-off.
@@
Table G: Ablation study for different rank settings of LoRA [22]. The model is trained on MARIO-LAION [12]. The number of parameters indicated on a million scale. We denote with bold for the selected rank configuration.
Impact of LoRA Rank. We examine the influence of various rank configurations for LoRA [22] and provide justification for the selected rank. We assess models trained with different rank values trained with the MARIO-10M dataset (Table G). We observe a consistent pattern of performance improvement with increasing rank values then saturates at $r=32$, thus we opt for a rank setting of $r=32$ as our final choice.
@@ Appendix C Additional Qualitative Results
Additional qualitative results for English texts. We provide additional qualitative results for English text image generation (Figure B). The model fine-tuned on AText is adopted for the English image generation. We use the MARIO-EVAL benchmark prompts [12] for the generation.
@@
Additional qualitative results for multi-lingual texts. We provide additional qualitative results for the multi-lingual, Latin text image generation (Figure C), and multi-lingual, non-Latin text image generation (Figure D). We use the model trained on the merged set of ICDAR2019 [14], and synthetic images of all the languages. Despite the deficiency of real samples of Russian, Greek, and Thai, our model shows the capability of generalizing to these languages.
@@
Additional qualitative results for logos. We present additional qualitative results of logo image generation. We first present the generation results of the logos that are included in the training set (Figure E). The results are generated using the logo benchmark prompts. The generated results confirm the model’s capability of generating logos in the desired location of the corresponding reference logo image, and that this reference extends beyond text renderings. Moreover, we present the generation results of the logos that are never seen during training (Figure F). The results further validate the effectiveness of the proposed method and the model’s ability to generalize across novel instances.
@@
Additional Text Image Editing Results. We provide additional text image Editing results in Figure G. The model successfully edits the specified region to include the intended words without modifying the rest of the region.
@@
Figure B: Additional results for English generation. The odd columns are the prompts, and the even columns are the generated images. Words enclosed with punctuation are the target keywords to be generated.
@@
Figure C: Additional qualitative results for multi-lingual, Latin text image generation. Images in the same row correspond to the same prompt templates. The first column denotes the prompts. The even column denotes the generated results, and the odd column denotes the reference image of text renderings.
@@
Figure D: Additional qualitative results for multi-lingual, non-Latin text image generation. Images in the same row correspond to the same prompt templates. The first column denotes the prompt template. The even column denotes the generated results, and the odd column denotes the reference image of text renderings.
@@
Figure E: Additional qualitative results for logo image generation. The reference logos utilized here are part of the training logo set. Images in the same row correspond to the same prompt templates. The first column denotes the prompt template. We replace “[KEYWORD]” with the name of the logo during the inference. The even columns denote the generated results, and the odd columns denote the reference logo images.
@@
Figure F: Additional qualitative results for logo image generation. The reference logos utilized here are not included in the training logo set and hence, are unseen during the model training. Images in the same row correspond to the same prompt templates. The first column denotes the prompt template. The even columns denote the generated results, and the odd columns denote the reference logo images.
@@
Figure G: Additional qualitative results for Text Image Editing. Regions denoted in yellow are masked during the inference. We denote the added text below each of the results.

### v2

@@ A.5 Text Image Editing Details
We elaborate on the extension of our model for text image editing. In line with [textdiffuser], we augment the input channel by incorporating additional encoded latent obtained from masked images. Specifically, we use VAE to encode an image containing masked regions and concatenate the encoded latent with the original model input. We use the same VAE used for reference image encoding to encode the masked image, and noise is not added to the resulting latent. Consequently, we obtain an augmented input $\tilde{\mathbf{x}}_{i}^{t}\in\mathcal{R}^{(3c+1)\times h\times w}$ as depicted below:
$\small\tilde{\mathbf{x}}_{i}^{t}=\text{concat}(\mathbf{z}_{i}^{t},\mathbf{s}_{i},\mathbf{m}_{i},\mathbf{z}_{i}^{\text{masked}}),$ (10)
@@ B.1 Ablation Study on Languages
ControNet 15.00 12.00 08.50 07.00 08.50 11.00 15.00 6.00 19.50 10.63 12.00 11.75 119.996
SD 0.00 0.50 0.0100 0.00 0.00 0.00 0.00 0.00 0.00 0.38 0.00 0.15 93.198
TextDiffuser 84.00 59.00 62.50 27.50 0.00 0.00 0.00 0.00 0.00 58.25 0.00 31.70 96.826
Ours 70.00 48.00 56.00 51.50 21.00 11.50 38.00 15.50 33.00 56.38 23.80 41.45 97.523
Ours 81.75 58.50 60.50 57.50 22.00 10.00 48.00 18.00 48.00 64.56 29.20 48.60 97.156
@@
Non-Latin Hindi 193.788 0.3125 21.00 37.54
Bengali 192.716 0.3135 11.50 17.60
Russian 188.567 0.2998 38.00 56.89
Thai 203.457 0.3015 15.50 27.59
Greek 192.215 0.3023 33.00 55.20
Average 97.523 0.3006 41.45 58.94
@@
Per language comparison.
We provide OCR Accuracy for each language and compare it with the existing methods in Table A. Remarkably, our model is shown to be the most accurate, as it surpasses existing methods across all average accuracy metrics, including the average of Latin, non-Latin, and the entire language set by a significant margin. Moreover, our method achieves high English accuracy comparable to that of TextDiffuser, which was tailored for English image generation. Although TextDiffuser shows high accuracy in some languages, such as English and German, it shows incapability in the rest of the language set, as it achieves 0 accuracy. In contrast to TextDiffuser, our method can generate text in all languages, and this demonstrates the multi-lingual capability of our model.
Latin v.s. non-Latin.
Our model exhibits higher OCR Accuracy scores in Latin languages such as English, Italian, German, and French (Tables A and B). It is noteworthy that for Russian, Thai, and Greek, our model is exclusively trained on synthetic images, yet it achieves comparable or superior FID and CLIP scores compared to the results for Latin languages. Given that the model was pretrained on a large-scale English dataset [textdiffuser], we speculate that this contributes as one factor to its superior performance in Latin languages.
Different levels of complexity among languages.
We note that different languages present different levels of complexity, with some being more challenging to generate accurately (Tables A and B). For instance, the size of the alphabet in English is 26, whereas in Thai, the size is 59. In addition, in languages such as Thai and Greek, diacritics are combined with other letters, and extra complexity is added. Some of the predictions that our model makes look similar to the actual word, but are evaluated to be incorrect due to the mis-generation of diacritics. Besides, the shapes of characters in non-Latin languages (e.g., Bengali), are notably more intricate than Latin characters, whose alphabet typically requires a greater number of strokes to form.
@@
0.5 0.5 7 38.104 0.3454 59.84
0.5 1.5 6 39.932 0.3448 59.25
0.5 2.5 5 42.192 0.3434 61.73
0.5 3.5 4 44.680 0.3419 62.51
0.5 4.5 3 49.827 0.3317 63.85
- 0 7.5 38.593 0.3454 58.26
Table D: Evaluation results on MARIO-EVAL with varying portions of scheduled guidance $\gamma$. $\gamma=0$ denotes the result of the CFG without the guidance scheduling.
@@ B.2 Ablation Study on CFG scheduling
$\small\begin{split}\hat{\epsilon}^{t}&=\underbrace{\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)}_{\text{Unconditional}}\\ &\underbrace{+\omega_{\text{prompt}}\Big[\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\mathbf{c},t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]}_{\text{Caption Guidance}}\\ &+\underbrace{\omega_{\text{ref}}\Big[\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\varnothing,t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]}_{\text{Reference Guidance}}\\ &+\underbrace{\omega_{\text{all}}\Big[\epsilon(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\mathbf{c},t)-\epsilon(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]}_{\text{All Guidance}},\end{split}$
@@
Impact of the scheduling speed.
We first study the impact of CFG scheduling by varying $\rho_{\text{speed}}$, the increase/decrease rate of the guidance scale, in Table D. We observe a rapid decrease in $\omega_{\text{ref}}$ with a rapid increase in $\omega_{\text{ref}}$ result in a decrease in the OCR Accuracy and improvements in FID and CLIP scores. As $\omega_{\text{ref}}$ decreases rapidly, the impact of the prompt gets higher and the reference condition becomes less influential in the generation process. As a consequence, non-target words within the prompt gain increased influence and affect the generation, resulting in the inclusion of non-target words in the output (Figure A).
Impact of different portions of scheduled guidance.
We examine the effect of different portions of scheduled guidance $\gamma$ within the total amount of guidance (Table D). The result of CFG without scheduling is also provided for comparison. We observe the general trend of a decrease in CLIP scores and an increase in OCR Accuracy as $\gamma$ increases. We speculate that the initial guidance scale of the reference condition plays an important role in establishing the initial layout of the text to be generated which, in turn, influences the effect of the prompt condition and consequently affects CLIP and FID.
@@ B.3 Ablation Study on Hyper Parameters
Impact of synthetic loss weight.
We first analyze the impact of the synthetic loss weight in Table F. We note that the synthetic loss weight mainly impacts the CLIP as the text prompts of the real dataset better describe the content of an image. We observe the model trained with a synthetic loss weight of 0.5 achieves a lower CLIP than the one trained with $1.0$, which validates the effectiveness of lowering the weight of reconstruction loss when synthetic images are provided.
Impact of recognition loss weight.
We analyze the impact of the recognition loss weight on accuracy and the quality of the generated results in Table F. We note that the OCR accuracy improves as the model is trained with increased recognition loss weight. In general, models with higher loss weight achieve better OCR Accuracy and a worse score on FID. We choose the recognition weight as $0.025$ for a proper trade-off.
@@
Table G: Ablation study for different rank settings of LoRA [lora]. The model is trained on MARIO-LAION [textdiffuser]. The number of parameters indicated on a million scale. We denote with bold for the selected rank configuration.
Impact of LoRA rank.
We examine the influence of various rank configurations for LoRA [lora] and provide justification for the selected rank. We assess models trained with different rank values trained with the MARIO-10M dataset (Table G). We observe a consistent pattern of performance improvement with increasing rank values, then saturates at $r=32$, thus we opt for a rank setting of $r=32$ as our final choice.
Type Acc $\uparrow$ F1 $\uparrow$ CLIP
Blank Reference 20.27 27.18 0.3695
Ours 42.87 48.86 0.3759
Table H: Impact of subject existence in logo generation.
Type Acc $\uparrow$ F1 $\uparrow$ CLIP
Blank Reference 1.88 2.71 0.2855
Ours 55.19 79,24 0.3468
Table I: Impact of subject existence in English generation.
[…]
SD-v1.5 0.03 0.3015
Ours-v1.5 51.40 0.3647
SD-v2.1 0.02 0.3221
Ours-v2.1 49.20 0.3685
Table J: Evaluation of different SD versions on the OpenLibraryEval500 subset.
[…]
English 58.26
English+Logo 51.64
Table K: Simultaneous training result for English generation.
B.4 Analysis on Reference Image.
Impact of reference image during inference.
To study the impact of the appearance of the reference image, we compare generation results using the reference image with and without subject rendering i.e., blank reference in Tables H and I. The results show that the generation result is greatly influenced by the appearance of the image. We also provide qualitative results in Figure B.
B.5 Ablation Study on SD Versions
To show generalization over different versions, we provide results on SD-v1.5 in Table J. Our method effectively improves on both versions.
B.6 Ablation Study on Simultaneous Task Training
Our method employs separate plugin modules, each acting as an expert for a specific task. When a task is prompted, its corresponding expert plugin is loaded. This modular design reduces interference between tasks, which often occurs when training with diverse objectives. Experimental results confirm this advantage: simultaneous task training leads to degraded performance (Table K).
Impact of subject location in reference.
We present a visual analysis of how the spatial position of the subject in the reference image affects the generated outputs. As illustrated in Figure D, the subject’s location in the reference image is consistently reflected in its position within the generated results.
Joint influence of prompt and reference.
We observe that both the input prompt and the reference image influence the generation results. To show this, we visualize the generation result with the same reference image, but with a different input prompt. As shown in Figure C, the generation results not only depend on the appearance of the subject in the reference image, but also on the input prompt.
Figure B: Generation result with blank reference image.
@@ Appendix C Additional Qualitative Results
We provide additional qualitative results for English text image generation (Figure E). The model fine-tuned on AText is adopted for the English image generation. We use the MARIO-EVAL benchmark prompts [textdiffuser] for the generation.
Additional qualitative results for multi-lingual texts.
@@
We provide additional qualitative results for the multi-lingual, Latin text image generation (Figure F), and multi-lingual, non-Latin text image generation (Figure G). We use the model trained on the merged set of ICDAR2019 [icdar2019], and synthetic images of all the languages. Despite the deficiency of real samples of Russian, Greek, and Thai, our model shows the capability of generalizing to these languages.
Additional qualitative results for logos.
@@
We present additional qualitative results of logo image generation. We first present the generation results of the logos that are included in the training set (Figure H). The results are generated using the logo benchmark prompts. The generated results confirm the model’s capability of generating logos in the desired location of the corresponding reference logo image, and that this reference extends beyond text renderings. Moreover, we present the generation results of the logos that are never seen during training (Figure I). The results further validate the effectiveness of the proposed method and the model’s ability to generalize across novel instances.
Comparison results with personalization methods.
We provide comparison results with personalization methods in Figure J.
Additional text image editing results.
@@
We provide additional text image Editing results in Figure K. The model successfully edits the specified region to include the intended words without modifying the rest of the region.
Figure C: Generation result with the same reference image. The images in the same row are generated with the same reference, but with different input prompts.
Figure D: Generation result with varying position.
@@
Figure E: Additional results for English generation. The odd columns are the prompts, and the even columns are the generated images. Words enclosed with punctuation are the target keywords to be generated.
@@
Figure F: Additional qualitative results for multi-lingual, Latin text image generation. Images in the same row correspond to the same prompt templates. The first column denotes the prompts. The even column denotes the generated results, and the odd column denotes the reference image of text renderings.
@@
Figure G: Additional qualitative results for multi-lingual, non-Latin text image generation. Images in the same row correspond to the same prompt templates. The first column denotes the prompt template. The even column denotes the generated results, and the odd column denotes the reference image of text renderings.
@@
Figure H: Additional qualitative results for logo image generation. The reference logos utilized here are part of the training logo set. Images in the same row correspond to the same prompt templates. The first column denotes the prompt template. We replace “[KEYWORD]” with the name of the logo during the inference. The even columns denote the generated results, and the odd columns denote the reference logo images.
@@
Figure I: Additional qualitative results for logo image generation. The reference logos utilized here are not included in the training logo set and hence, are unseen during the model training. Images in the same row correspond to the same prompt templates. The first column denotes the prompt template. The even columns denote the generated results, and the odd columns denote the reference logo images.
Figure J: Comparison results with personalization methods.
@@
Figure K: Additional qualitative results for Text Image Editing. Regions denoted in yellow are masked during the inference. We denote the added text below each of the results.